The random variable x has the discrete probability distribution shown here:

x -2 -1 0 1 2

P(x) 0.10 0.15 0.40 0.30 0.05


Find P(x ≤ 0)

b. Find P(x > -1).

c. Find P(-1 ≤x≤ 1) .

d. Find P(x <2) .

e. Find P(-1 < x < 2) .

f. Find P(x < 1) .


1
Expert's answer
2021-08-16T09:55:16-0400


a. P(x ≤ 0) = 0.10 + 0.15 +0.40 = 0.65

b. P(x > -1) = 0.40+0.30+0.05 = 0.75

c. P(-1 ≤x≤ 1) = 0.15+0.40+0.30=0.85

d. P(x <2) = 0.10 + 0.15 + 0.40+ 0.30 = 0.95

e. P(-1 < x < 2) = 0.40+0.30 = 0.70

f. P(x < 1) = 0.10 + 0.15 + 0.40 = 0.65